Randomly select 24 component and accept the whole batch if there are fewer then 3 defective. If a particular shipment of thousands of components actually has 4% rate of defects. What is the probability that this whole shipment will be accepted?
1
Expert's answer
2020-04-10T17:54:29-0400

Let "X=" the number of defective components: "X\sim Bin(n,p)"


"P(X=x)=\binom{n}{x}p^x(1-p)^{n-x}"

Given that "n=24,p=0.04"

The probability that this whole shipment will be accepted is


"P(X<3)=P(X=0)+P(X=1)+P(X=2)="

"=\binom{24}{0}0.04^0(1-0.04)^{24-0}+\binom{24}{1}0.04^1(1-0.04)^{24-1}+"

"+\binom{24}{2}0.04^2(1-0.04)^{24-2}\approx"

"\approx0.375413+0.375413+0.179886\approx0.9307"